The horizontal range of a projectile is R and the maximum height attained by it is H. A strong wind now begins to blow in the direction of motion of the projectile, giving it a constant horizontal acceleration= g/2. Under the same conditions of projectile, find the horizontal range of the projectile

Grade:11

6 Answers

mohit vashishtha
8 years ago
new range will be R+2H
APPROVE IF USEFUL
AP
8 years ago
How? Will you please explain?
Hasan Naqvi
8 years ago
Time taken by projective to reach ground = 2 * (2H/g)1/2
S= ut + ½ at2
= R + ½ * g/2 *2* ((2H/g)1/2)2
=R + g * H/g
=R+ H 
Nicho priyatham
8 years ago
R+2H 
we know that in 1st  case  H=(u2sin2a)/2g   R=(u2sin2a)/g
now wright the equation of motion in second case
s=((ucosa)t-(1/2)(g/2)t2 )i-((usina)t-(1/2)gt2 )j
when partical lands on ground coffeicant of j =0 so t=(2usina)/g subatute above equation in cofficant of i to  get new range
so we get (u2sin2a)/g +(u2sin2a)/g
=R+2H
 
kritika sharma
6 years ago
so its total time of flight and Hmax will be same T=2Vy/g & Hmax=Vy^2/2g
because of the wind horizontal acc. will be g so range will change
Rnew=VxT+1/2 gT^2 =2VxVy/g + 2Vy^2/g
so, 2VxVy/g is normal range & 2Vy^2/g=4H
so ans. is R+4H.
 
Raahul
5 years ago
😋😋😋😋😋🖐🖐🖐😊😋😋😋😋😊🖐🖐😋🖐😋🖐😋🖐😊😊😋😔😊😔😋😔😊😋😔😋🖐😘😆😁😍😘🖐😆🤣😭😗🖐😆😃🤗😗😭😃😶🤣😁😉🖐👌👍🖐😀🤗😁😁😭😗😊😋😀🤗😁